xXMarcelieXx:

help please.. i need help with number 7

xXMarcelieXx:

1 attachment
xXMarcelieXx:

i did part a

1 attachment
xXMarcelieXx:

part b which is a true stmt

1 attachment
xXMarcelieXx:

i did a set up for the proof but i got stuck on the backwards direction

1 attachment
Hero:

So you need help solving for a?

xXMarcelieXx:

i need help with b since i was doing the bi conditional proof

xXMarcelieXx:

i got stuck on the backwards direction

Hero:

Yes, but in b it says to "solve for a" do you see that?

Hero:

Should I circle it for you?

xXMarcelieXx:

oh wait.. it says part a lol umm how do i determine if they are true?

Hero:

You seem confused

xXMarcelieXx:

yeah i am :/

Hero:

You should identify exactly where you need help.

xXMarcelieXx:

hmm, well i need help with b. Where do i start? i think my work is wrong

Hero:

I don't know what "b" means. I don't see a "b"

Hero:

I already asked you if you needed help solving for a but it seems you did not recognize what I was talking about

xXMarcelieXx:

1 attachment
Hero:

Okay, but I was looking at your "proofs". There is a part where it says "solve for a" and you did not solve for a. Why did you not solve for a?

xXMarcelieXx:

i didnt solve it since i was lost.

Hero:

You were lost because you could not solve the equation for a?

xXMarcelieXx:

well i think i solved it which i created the bi conditional statement for it

Hero:

I'm referring to the Backwards Direction P2

xXMarcelieXx:

oh okay, sorry i got lost. Yes, i could solve for which i'll get it into factored form

gracie101507:

i dont know what question you are talking about send it one more time plz

Hero:

He already posted it twice

katwhirles:

the picture is at the top and question 7

katwhirles:

i am sorry i cant help

gracie101507:

I am sorry i can¨t help so sorry

katwhirles:

becsuse i really dont know the answer

katwhirles:

because

gracie101507:

i dont really now the answer i am in 8th grade

xXMarcelieXx:

im completely lost tbh lol i solved for a already for i ended up factoring it , which i think its wrong

Hero:

@xXMarcelieXx to solve for \(a\) you have to write it in vertex form by completing the square

gracie101507:

oh that sounds about right

gracie101507:

did you figure out part A. and part B

xXMarcelieXx:

oh wait i think got it one sec

xXMarcelieXx:

1 attachment
Hero:

Yes correct

xXMarcelieXx:

okay so then, we would square each side and we up having sqroots but isnt that hard to solve?

gracie101507:

i really dont know

Hero:

In forward direction, when you solved for a, what did you do with the expression after that?

xXMarcelieXx:

i plugged in a into the conclusion and it was easier to solve

Hero:

So why not do that for the backwards direction as well?

xXMarcelieXx:

i did this but i cant get it into the form of 8p though

1 attachment
Hero:

One moment

xXMarcelieXx:

okie

gracie101507:

what grade are you in xXMarcelieXx

gracie101507:

it is kinda off topic but i kinda what to know

Hero:

I have something that might work.

Hero:

I don't know where @xXMarcelieXx went to though

Hero:

$$ \begin{array}{l} a^{2} +4a-4=8r\\ a^{2} -8a+4a-4=8r\ -\ 8a\\ a^{2} -4a-4=8r-8a\\ a^{2} -4a+4-4=8r-8a+4\\ a^{2} -4a+4=8r-8a\\ ( a-2)^{2} =( 8r-8a)\left(\dfrac{8r+8a}{8r+8a}\right)\\ ( a\ -\ 2)^{2} =\dfrac{( 8r-8a)( 8r+8a)}{8r+8a}\\ ( a\ -\ 2)^{2} =\dfrac{64r^{2} -64a^{2}}{8r+8a}\\ ( a\ -\ 2)^{2} =\dfrac{64( r+a)( r-8)}{8( r+a)}\\ ( a-2)^{2} =\dfrac{64}{8}( r-8)\\ a-2=\sqrt{\dfrac{64}{8}( r-8)}\\ a-2=\sqrt{64}\sqrt{\dfrac{r-8}{8}}\\ a\ -2=\ 8\sqrt{\dfrac{r-8}{8}} \end{array}$$

Hero:

Somehow we'd have to show that root expression is equal to \(p\)

Hero:

And if it can't be proven then maybe the original statement is false

Hero:

Also I just realized something is off with your work

xXMarcelieXx:

oh, where is it off?

Hero:

Basically, what I'm trying to figure out is, how did you come up with your table of values?

Hero:

You're using two variables I think when you should just be using \(a\) as the variable

Hero:

That's how you've always done these unless the original statement had two variables which it does not

Hero:

Remember, when we are doing these normally, the other variable isn't introduced until later. We should keep that same consistency here

Hero:

I'm referring to your table of true/false values btw

xXMarcelieXx:

right, i messed on my algebra since i didnt complete the square right. i thought we always used two variables for this stuff?

Hero:

Negative.

Hero:

Not for verifying the original statement

Hero:

The original statement only has one variable to use for verifying it which is variable \(a\)

Hero:

Do you get what I'm saying?

Hero:

As I said, when doing a proof, you only introduce another variable AFTER you have done verifying with tables

Hero:

You're still more confused than you should be

xXMarcelieXx:

oh okay, i think i get what you are saying

xXMarcelieXx:

how did you get this second step ?

1 attachment
Hero:

I obviously subtracted 8a from both sides

Hero:

Which you are allowed to do

Hero:

Actually I realize the statement should be \((a - 2)^2 = 8r - 8a + 8\) \((a - 2)^2 = 8(r - a + 1)\)

Hero:

One moment

xXMarcelieXx:

okie

Hero:

\(\begin{array}{l} a^{2} +4a-4=8r\\ a^{2} -8a+4a-4=8r\ -\ 8a\\ a^{2} -4a-4=8r-8a\\ a^{2} -4a+4-4=8r-8a+4\\ a^{2} -4a+4=8r-8a\ +\ 8\\ ( a\ -\ 2)^{2} =8r\ +\ 8\ -\ 8a\left(\dfrac{8r+8+8a)}{8r+8+8a}\right)\\ ( a\ -\ 2)^{2} =\dfrac{( 8r\ +\ 8\ -8a)( 8r+8+8a)}{8r+8+8a}\\ ( a\ -\ 2)^{2} =\dfrac{(( 8r\ +\ 8) -8a)(( 8r+8) +8a)}{8r+8+8a}\\ ( a\ -\ 2)^{2} =\dfrac{( 8r+8)^{2} -( 8a)^{2}}{8r+8+8a}\\ ( a\ -\ 2)^{2} =\dfrac{( 8( r+1))^{2} -( 8a)^{2}}{8r+8+8a}\\ ( a\ -\ 2)^{2} =\dfrac{64( r+1)^{2} -64a^{2}}{8( r\ +\ 1+a)}\\ ( a-2)^{2} =\dfrac{64\left(( r+1)^{2} -a^{2}\right)}{8( r+1+a)}\\ ( a-2)^{2} =\dfrac{64}{8} \cdot \dfrac{( r+1+a)( r+1-a)}{( r+1+a)}\\ ( a-2)^{2} =\dfrac{64}{8}( r+1-a)\\ a-2=\sqrt{\dfrac{64}{8}( r+1-a)}\\ a-2=\sqrt{64}\sqrt{\dfrac{r+1-a}{8}}\\ a-2=8\sqrt{\dfrac{r+1-a}{8}} \end{array}\)

Hero:

So you'd have to prove that \(p\) is equal to that root expression

Hero:

But the expression has an a in it which is not good

Hero:

I'm leaning towards the statement is false

xXMarcelieXx:

ohhh, so then does that mean that the whole statement is false?

Hero:

Hasn't exactly been verified whether true or false

Hero:

But what you can do is select an \(a\), then see if you can find an \(r\) such that \(p\) is an integer.

Hero:

The root expression must be equal to \(p\)

Hero:

Hang on a min ...

Hero:

Basically the root expression is equal to \(p\)

Hero:

So it is proven

Hero:

I did it by accident

Hero:

\(p = \sqrt{\dfrac{r + 1 -a}{8}}\)

Hero:

Not sure if your teacher will accept it though since \(a\) is included in the expression for \(p\)

Hero:

But either way, its a formula that takes values \(a\) and \(r\) and returns a \(p\)

Hero:

where \(a\) , \(r\) and \(p\) are integers that satisfy the given condition

Hero:

For example \(a = 10\) and \(r = 17\) produces \(p = 1\)

Hero:

You can produce infinite of these values

Hero:

You can always find an \(a\) and and \(r\) that produces a \(p\)

Hero:

Do you follow @xXMarcelieXx

xXMarcelieXx:

oh okay, got it :)

Hero:

@xXMarcelieXx have you submitted this to your teacher yet?

xXMarcelieXx:

xXMarcelieXx:

i got confused what she wrote after the or stuff

Hero:

I'd have to review mods in order to continue helping you with them. I don't know how to manipulate them like that.

xXMarcelieXx:

oh okay, i remember i did the the first part of (p->Q) with mods but the second part was too much